Fermi National Laboratory


Accelerator Update

Friday July 23
The day shift began with machine experts conducting studies, and with beam to MiniBooNE and the Meson MCenter and MTest experiments.

Machine studies ended around 1:15 PM.

At 3:42 PM, the MiniBooNE horn tripped off for the first of several times. Operations called in an expert at 4:18 PM.

Operations put the Tevatron (TeV) through the first of several dry squeezes at 6:09 PM.

Operations reported at 6:28 PM that the MiniBooNE horn was operationally normal.

At 8:22 PM, Operations couldn't turn on the Tevatron Electron Lens (TEL). Operators called in an expert who had the device on twenty minutes later.

At 10:19 PM, a Main Injector (MI) RF station (MIRF8) tripped off for the first of several times. It reset each time okay.

Operations loaded a 12x0 study-store at 10:27 PM.

MTest Experimenters asked at 11:20 PM that their beam be disabled.

Saturday July 24
The midnight shift began with TeV experts conducting studies and with MiniBooNE taking beam.

Operations terminated the study-store at 12:14 AM.

Operations put the TeV into shot setup at 12:21 AM.

At 1:57 AM, MIRF8 began tripping off again. An MI expert tried to tune coalescing from MI-60, but couldn't. She reconfigured the MI RF system to run without MIRF8.

Operations established store 3671 at 3:47 AM with an initial luminosity of 28.88E30. Pbar resumed stacking. Operator began investigating the cause of the low luminosity.

MIRF8 was turned back on at 4:11 AM.

Operation established beam to MCenter at 5:02 AM.

Operations put stacking on hold at 5:53 AM. Operators, with the help of a TeV expert, traced the low luminosity to a TeV separator being at the wrong voltage. Operators set the separator to the proper voltage and then adjusted the collimators. Operations reported at 6:05 AM that the luminosity was at 62.18E30.

Operations began setting up MTest to take beam at 9:43 AM and then disabled beam to MTest at 3:34 PM.

Operations established beam to MCenter at 7:19 PM

At 8:57 PM, a Booster power supply (EAPS) tripped off due to a RF station (BRF4) problem. BRF4 wouldn't reset. Operations reconfigured the RF system at 9:16 PM to operate without this station at a lower intensity. An expert was called in who had the station back on and in the system by 10:17 PM.

BRF1 tripped off at 9:22 PM, for an unknown reason, for the first of three times in thirty minutes. It reset okay each time.

Pbar resumed stacking at 10:20 PM.

A Linac RF station (KRF5) tripped off at 10:31 PM. It reset okay.

BRF4 tripped off at 10:42 PM. It reset okay.

KRF5 tripped again at 11:44 PM. It reset okay.

Sunday July 25
The midnight shift began with Operations monitoring store 3671, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

MIRF11 tripped of at 1 AM. It reset okay.

LRF2 tripped off at 4:20 AM. It required operator intervention to reset.

Operations terminated store 3671 at 5 AM.

Operations put the TeV into shot setup at 6:07 AM. Operations disabled the Switchyard (SY) beam for shot setup.

At 8:02 AM, MIRF8 began tripping again. It was left off and Operations reconfigured MI to run with seventeen RF stations.

Operations established store 3674 at 9:07 AM with an initial luminosity of 82.2E30. Pbar resumed stacking.

MCenter went down for the day. Operations established beam to MTest at 10:28 AM.

At 10:45 AM, BRF1 tripped off, and was reset more times than there was room here to record over the next eight hours.

Operations established beam to MCenter at 11:20 AM and disabled beam to MTest at 12:03 PM.

At 3:46 PM, BRF11 tripped repeatedly. It reset okay.

At 4:45 PM, BRF17 began tripping on a power supply overload. Operators could not get it to stay on. An expert was called in. Operations disabled the MiniBooNE beam while they investigated. At 6:30 PM, the RF expert took BRF17 out of the RF system. Beam to MiniBooNE resumed at 6:51 PM.

The RF expert investigated the BRF1 problems and had it back online at 6:52 PM.

MTest asked to have their beam disabled at 7:23 PM.

At 7:23 PM, the MiniBooNE horn tripped off. Operators reset the device, but it tripped off again at 7:44 PM on a ground fault indication. Operators called in an expert.

BRF14 tripped off at 8:13 PM. It reset okay.

At 8:42 PM, Operations reported that they had to slow stacking in order to control emittances.

Operations turned on MIRF8 and brought back into the RF system at 10 PM.

A Booster device (CEPSNE2) tripped off at 10:11 PM for the first of four times in three hours.

Monday July 26
The midnight shift began with Operations monitoring store 3674, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam

MTest asked to have their beam disabled at 12:36 AM.

At 2:02 AM, the MiniBooNE horn tripped off. A MI expert said to lower the horn voltage.

At 3:08 AM, the MiniBooNE horn tripped off again. Operators had to lower the voltage again to get it to stay on.

At 4:58 AM, a Linac quadrupole power supply (QPS107) went out of tolerance. Operators adjusted it back to its nominal reading.

At 5:51 AM, CEPNE2 tripped off for the first of three times in about twenty-five minutes.

Plans for Monday
TeV experts will be allowed to conduct an end-of-store study. Operations will terminate store 3674 at 1 PM. The Collider Detector at Fermilab and D-Zero will be allowed an access period before the next store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/26/2004   email Fermilab